Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability, 3 dice simultaneously

Expert replies
by euro » Wed Sep 29, 2010 7:18 am
(Q) What is the probability of getting a sum of 8 or 14 when rolling 3 dice simultaneously?

a) 1/6
b) 1/4
c) 1/2
d) 21/216
e) 32/216

The official answer is [spoiler](a) [/spoiler]

Can somebody explain?
Join the discussion
Source: — Problem Solving |

by kmittal82 » Wed Sep 29, 2010 8:09 am
Wow, tough one...

Lets look at it case by case

#1) To get 8

8 = :

1+1+6
2+2+4
3+3+2
1+2+5
1+3+4


Since order won't matter, any combination of those 3 numbers will give you a sum of 8. Each combination can further have 6 arrangements


#2) to get 14
14 = :
6+6+2
6+5+3
6+4+4
5+5+4

Since order won't matter, any combination of those 3 numbers will give you a sum of 14. Each combination can further have 6 arrangements

Total outcomes on a roll = 6x6x6 = 216

Fav cases for getting 8 = 5x6/216 = 30/216
Fav cases for getting 14 = 4x6/216 = 24/216

Total probability = 54/216

Unfortunately though, its in neither of the answers, so I know Ive gone wrong somewhere... but where? :)
Join the discussion

by diebeatsthegmat » Wed Sep 29, 2010 9:30 am
euro wrote:(Q) What is the probability of getting a sum of 8 or 14 when rolling 3 dice simultaneously?

a) 1/6
b) 1/4
c) 1/2
d) 21/216
e) 32/216

The official answer is [spoiler](a) [/spoiler]

Can somebody explain?
for sum of 8 we will have
116 ( 161,611) thus there is 3 combination*2 =6 ( we have to multiply with 2 because the 3 dice are all different)
152 ,125, 215,251,512,521 thus there are 6 combination
the same with 143 we have 6 combination
total=18 ways thus probability=18/6*6*6
for sume of 14, we do the same above
266 , 635,464 and also have 18 ways
the probability is 18/6*6*6
total probability is 18/216+18/216=36/216=1/6
hope it helps
Join the discussion

by Stuart@KaplanGMAT » Wed Sep 29, 2010 9:56 am
kmittal82 wrote:Wow, tough one...

Lets look at it case by case

#1) To get 8

8 = :

1+1+6
2+2+4
3+3+2
1+2+5
1+3+4


Since order won't matter, any combination of those 3 numbers will give you a sum of 8. Each combination can further have 6 arrangements
The bolded statement is your error.

For three different digits, such as 1+3+4, there are indeed 6 different arrangements. However, for the cases that have only 2 distinct digits, there are only 3 different arrangements.

So:

1/1/6 - 3 arrangements
2/2/4 - 3 arrangements
3/3/2 - 3 arrangements
1/2/5 - 6 arrangements
1/3/4 - 6 arrangements

Total - 21 arrangements

Similarly for 14:

6/6/2 - 3 arrangements
6/4/4 - 3 arrangements
5/5/4 - 3 arrangements
6/5/3 - 6 arrangements

Total - 15 arrangements

Total possibilities is 6^3 = 216

Prob = # of desired outcomes / total # of possibilities

Prob = (21 + 15)/216 = 36/216 = 1/6
Image

Stuart Kovinsky | Kaplan GMAT Faculty | Toronto

Kaplan Exclusive: The Official Test Day Experience | Ready to Take a Free Practice Test? | Kaplan/Beat the GMAT Member Discount
BTG100 for $100 off a full course
Join the discussion

by kmittal82 » Wed Sep 29, 2010 11:30 am
Thanks for the explanation everyone :)
Join the discussion

by diebeatsthegmat » Wed Sep 29, 2010 4:59 pm
Stuart Kovinsky wrote:
kmittal82 wrote:Wow, tough one...

Lets look at it case by case

#1) To get 8

8 = :

1+1+6
2+2+4
3+3+2
1+2+5
1+3+4


Since order won't matter, any combination of those 3 numbers will give you a sum of 8. Each combination can further have 6 arrangements
thanks for the detail expanation but i thought the dices are different , of course it have the same numbers but its from different dice

The bolded statement is your error.

For three different digits, such as 1+3+4, there are indeed 6 different arrangements. However, for the cases that have only 2 distinct digits, there are only 3 different arrangements.

So:

1/1/6 - 3 arrangements
2/2/4 - 3 arrangements
3/3/2 - 3 arrangements
1/2/5 - 6 arrangements
1/3/4 - 6 arrangements

Total - 21 arrangements

Similarly for 14:

6/6/2 - 3 arrangements
6/4/4 - 3 arrangements
5/5/4 - 3 arrangements
6/5/3 - 6 arrangements

Total - 15 arrangements

Total possibilities is 6^3 = 216

Prob = # of desired outcomes / total # of possibilities

Prob = (21 + 15)/216 = 36/216 = 1/6
Join the discussion